OpenStudy (konradzuse):

Find a geometric power series for 1/(7-x) centered at 0

OpenStudy (konradzuse):

\[\frac{1}{(7(1-\frac{x}{7})}\]

OpenStudy (konradzuse):

\[\frac{1}{7} \sum_{n=0}^{\infty} (\frac{x}{7})^n\]

OpenStudy (konradzuse):

\[\sum_{n=0}^{\infty} (\frac{x^n}{7^{n+1}})\]

OpenStudy (saifoo.khan):

@SmoothMath , help?

OpenStudy (konradzuse):

:(?

OpenStudy (konradzuse):

:)?

OpenStudy (experimentx):



OpenStudy (experimentx):